A total of $6000 is invested: part at 9% and the remainder at 14%. How much is invested at each rate if the annual interest is $740?
Answer by mananth(16946)

Part I 9.00% per annum ------------- Amount invested =x
Part II 14.00% per annum ------------ Amount invested = y
6000
Interest----- 740.00

Part I 9.00% per annum ---x
Part II 14.00% per annum ---y
Total investment
x + 1 y= 6000 -------------1
Interest on both investments
9.00% x + 14.00% y= 740
Multiply by 100
9 x + 14 y= 74000.00 --------2
Multiply (1) by -9
we get
-9 x -9 y= -54000.00
Add this to (2)
0 x 5 y= 20000
divide by 5
y = 4000
Part I 9.00% $ 2000
Part II 14.00% $ 4000

CHECK
2000 --------- 9.00% ------- 180.00
4000 ------------- 14.00% ------- 560.00
Total -------------------- 740.00

m.ananth@hotmail.ca